# Preconditioning without a preconditioner: faster ridge-regression and Gaussian sampling with randomized block Krylov subspace methods

### v1

@@ 3.2 Computational costs
We emphasize that after the block Lanczos algorithm has been run, the cost to compute the block-CG iterates for multiple values of $\mu$ is relatively small. This is in contrast to Nyström PCG, which is is not particularly well suited for solving Equation 1 for multiple values of $\mu$; while the Nyström approximation can be reused, PCG must be re-run and new products with $\mathbf{A}$ computed.
3.3 Comparison with past work
The block-CG algorithm is widely used to solve linear systems with multiple right hand sides, and in such settings working over the block Krylov subspace is often advantageous compared to working over the individual Krylov subspaces (O’Leary, 1980; Meurant, 1984; Feng et al., 1995; Birk & Frommer, 2013). To the best of our knowledge, ours is the first work to consider a randomized variant of the block-CG algorithm as well as the first to prove that block-CG can outperform standard CG, even with the goal of solving a single linear system.
More broadly, a number of recent works in randomized numerical linear algebra (Chen & Hallman, 2023; Tropp & Webber, 2023; Meyer et al., 2024) have made use of various nestedness properties of block-KSMs similar in flavor to Theorem 4.1 in order to obtain more efficient algorithms.
4 Main analysis
In this section we give our theoretical analysis of augmented block-CG. In light of Theorem 3.2, our main strategy for deriving bounds for block-CG is simply to derive bounds for Nyström PCG. We do not attempt to optimize constants, opting instead for simple arguments and clean theorem statements.
4.1 Implicit preconditioning with block KSMs
We begin with a key observation.
Theorem 4.1.
Suppose $\mathbf{P}_{\mu}=(\mathbf{I}+\mathbf{X})^{-1}$ for some $\mathbf{X}$ with $\operatorname{range}(\mathbf{X})\subseteq\mathcal{K}_{s+1}(\mathbf{A}_{\mu},\mathbf{\Omega})$. Then,
A full proof is contained in Section B.1, but the basic idea is simple. By definition, $\mathcal{K}_{t}(\mathbf{A}_{\mu},\mathbf{b};\mathbf{P}_{\mu})$ consists linear combinations of the vectors
$(\mathbf{P}_{\mu}^{-1}\mathbf{A})^{k}\mathbf{P}_{\mu}^{-1}\mathbf{b}=((\mathbf{I}+\mathbf{X})\mathbf{A})^{k}(\mathbf{I}+\mathbf{X})\mathbf{b},$ (19)
where $k=0,1,\ldots,t-1$. And each $((\mathbf{I}+\mathbf{X})\mathbf{A})^{k}(\mathbf{I}+\mathbf{X})\mathbf{b}$ can be expressed as linear combination of vectors which live in the specificed space.
Proof of Theorem 3.2.
Clearly $\mathcal{K}_{t}(\mathbf{A},\mathbf{b})+\mathcal{K}_{t+s}(\mathbf{A},\mathbf{\Omega})\subseteq\mathcal{K}_{t+s}(\mathbf{A},[\mathbf{b}~{}\mathbf{\Omega}])$, so the result follows immediately from Theorem 4.1 and the optimality of block-CG and preconditioned-CG. ∎
4.2 Bounds for Nyström preconditioning
We begin by recalling a error guarantee for Nyström low-rank approximation, which compares the Nyström error $\|\mathbf{A}-\mathbf{A}\langle\mathbf{K}_{s}\rangle\|$ to the error $\|\mathbf{A}-[\![\mathbf{A}]\!]_{r}\|=\lambda_{r+1}$ of the best possible rank-$r$ approximation to $\mathbf{A}$.
Theorem 4.2 (Theorem 9.1 in (Tropp & Webber, 2023)).
Suppose $\mathbf{\Omega}\in\mathbb{R}^{d\times(r+p)}$ is a random Gaussian matrix and define
$\mathbf{K}_{s}:=[\mathbf{\Omega}~{}\mathbf{A}\mathbf{\Omega}~{}\cdots~{}\mathbf{A}^{s-1}\mathbf{\Omega}].$ (20)
Then, if $p\geq 2$,
$\log\Biggl{(}\frac{\operatorname{\mathbb{E}}\|\mathbf{A}-\mathbf{A}\langle\mathbf{K}_{s}\rangle\|^{2}}{\lambda_{r+1}^{2}}\Bigg{)}$
$\hskip 30.00005pt\leq\frac{1}{8(s-\frac{3}{2})^{2}}\log\Biggl{(}4+\frac{4r}{p-1}\sum_{i>r}\frac{\lambda_{i}^{2}}{\lambda_{r+1}^{2}}\Bigg{)}^{2}.$
The other main result that we need is a deterministic bound for the condition number of the Nyström preconditioned system, similar to Proposition 5.3 in (Frangella et al., 2023). The proof, which is contained in Appendix B, follows the exact same approach, although we state the bound for arbitrary $\theta>0$ while (Frangella et al., 2023) states it for one particular choice.
Theorem 4.3.
@@ 4.2 Bounds for Nyström preconditioning
Let $\mathbf{P}_{\mu}$ be the Nyström preconditioner Equation 8 corresponding to the Nyström approximation $\mathbf{A}\langle\mathbf{K}\rangle$ for any $\mathbf{K}$ and shift parameter $\theta\geq 0$. Then
$\kappa(\mathbf{P}_{\mu}^{-1/2}\mathbf{A}_{\mu}\mathbf{P}_{\mu}^{-1/2})$
@@
$\hskip 20.00003pt\leq\big{(}\theta+\mu+\|\mathbf{A}-\mathbf{A}\langle\mathbf{K}\rangle\|\big{)}\left(\frac{1}{\theta+\mu}+\frac{1}{\lambda_{d}+\mu}\right).$
To prove our main convergence guarantee for our augmented block-CG, we derive a new bound for Nyström preconditioning.
Theorem 4.4.
Let $\mathbf{\Omega}\in\mathbb{R}^{d\times(r+p)}$ be a random Gaussian matrix and define
$\mathbf{K}_{s}:=[\mathbf{\Omega}~{}\mathbf{A}\mathbf{\Omega}~{}\cdots\mathbf{A}^{s-1}\mathbf{\Omega}].$ (21)
Let $\mathbf{P}_{\mu}$ be the Nyström preconditioner Equation 8 corresponding to the Nyström approximation $\mathbf{A}\langle\mathbf{K}_{s}\rangle$ for any shift parameter $\theta\in[\lambda_{d},\lambda_{r+1}]$.
Suppose $p\geq 2$ and
$s\geq\frac{3}{2}+\frac{1}{2}\log\left(4+\frac{4r}{p-1}\sum_{i>r}\frac{\lambda_{i}^{2}}{\lambda_{r+1}^{2}}\right).$
Then, with probability at least $99/100$,
$\bigg{\{}\forall\mu:\kappa(\mathbf{P}_{\mu}^{-1/2}\mathbf{A}_{\mu}\mathbf{P}_{\mu}^{-1/2})\leq 28\frac{\lambda_{r+1}(\mathbf{A})+\mu}{\lambda_{d}+\mu}\bigg{\}}.$
Theorem 4.4, which we prove in Section B.2, follows from Theorems 4.2 and 4.3 and a simple application of Markov’s inequality. Our bound for Nyström PCG applies even when $\mu=0$, whereas the bounds in (Frangella et al., 2023) require $\mu$ to be sufficiently large relative to the sketching dimension $r+p$ and consider only $s=1$.
Theorem 4.4 may be hard for users of Nyström PCG to instantiate, since it is stated in terms of the (unknown) eigenvalues of $\mathbf{A}$. If $p=2$ then
$s\geq 3+\log(d)$ (22)
is sufficient to guarantee the condition on $s$ [^5] Indeed, $r\leq n$, $\lambda_{i}/\lambda_{r+1}\leq 1$ for $i>r$, and $\log(8)+2\log(d)=\log(d^{2})+\log(2)+\log(4)\geq\log(1+d^{2})+\log(4)=\log(4+4d^{2})$, and $33/2+\log(8)/2<$ 3. However, the user must still guarantee $\theta\in[\lambda_{d},\lambda_{r+1}]$, and different choices may lead to better or worse convergence. Theorem 3.2 implies that choosing $\theta$ is uneccessary for our augmented block-CG method.
Figure 2: Error versus matrix-loads for block-CG (Refer to caption), CG (Refer to caption), and Nyström PCG with $s=1$ (Refer to caption) and $s=3$ (Refer to caption) on several test problems.
4.3 Bounds for augmented block-CG
We can now state our main convergence guarantee for augmented block-CG. This result is a simple consequence of our condition number bound Theorem 4.4 and the error guarantee Corollary 2.3 for PCG.
@@ 4.3 Bounds for augmented block-CG
$s\geq\frac{5}{2}+\frac{1}{4}\log\left(4+\frac{4r}{p-1}\sum_{i>r}\frac{\lambda_{i}^{2}}{\lambda_{r+1}^{2}}\right),$
and
$\varepsilon_{t}(\mu):=2\exp\left(-\frac{t-s}{\sqrt{(\lambda_{r+1}+\mu)/(\lambda_{d}+\mu)}}\right).$
@@
$\Bigg{\{}\forall\mu:\frac{\|\mathbf{A}_{\mu}^{-1}\mathbf{b}-\mathsf{bcg}_{t}^{(1)}(\mu)\|_{\mathbf{A}_{\mu}}}{\|\mathbf{A}_{\mu}^{-1}\mathbf{b}\|_{\mathbf{A}_{\mu}}}\leq\varepsilon_{t}(\mu)\Bigg{\}}.$
Proof of Corollary 3.4.
Set $p=2$ and $s\geq 3+\log(d)$. By Equation 22 we can instantiate Theorem 4.5. ∎
4.4 Can increasing the Nyström depth parameter help?
As noted in Remark 3.3, our main result Theorem 3.2 allows us to compare augmented block-CG to Nyström PCG for a range of low-rank approximation depths $s$. Unfortuantely, standard bounds for randomized KSMs for low-rank approximation are for “large-block” methods which use a block size $\ell$ comparable to the target rank $r$ (Halko et al., 2011; Musco & Musco, 2015; Tropp & Webber, 2023). As a result, our bounds for Nyström preconditioning (e.g. Theorems 4.4 and 4.5) do not improve substantially as $s$ increases, and so there is no benefit in “optimizing” over $s$.
Note that $\mathbf{A}\langle\mathbf{K}_{s}\rangle$ as defined in Equation 11 can have rank as large as $\ell s$. Therefore, we might hope that it is competitive with the best rank-$\ell s$ approximation to $\mathbf{A}$; i.e. that
@@ 4.4 Can increasing the Nyström depth parameter help?
$\|\mathbf{A}-\mathbf{A}\langle\mathbf{K}_{s}\rangle\|\approx\lambda_{\ell s+1}.$ (23)
While this is not always possible (e.g. if $\mathbf{A}$ had many repeated top eigenvalues), recent work (Meyer et al., 2024) proves that for rank-$r$ approximation, certain “small-block” KSMs (e.g. with $\ell=O(1)$) can perform nearly as well as “large-block” KSMs (e.g. with $\ell=O(r)$) in terms of matrix-products under mild assumptions on eigenvalue gaps. Extending the ideas of (Meyer et al., 2024) to the Nyström approximation Equation 11 would immediately provide better results for Nyström PCG and hence our augmented block-CG.
5 Numerical experiments
We use the block-Lanczos algorithm with full-reorthogonalization in order to implement block-CG and the block-Lanczos square root iterates. Implementation details are described in Appendix A. The $s$ matrix-loads required to build the Nyström preconditioner are accounted for in the plots. We use the same random Gaussian matrix $\mathbf{\Omega}\in\mathbb{R}^{d\times\ell}$ for Nyström PCG and block-CG. More numerical experiments, including without full-reorthogonalization, are included in Appendix E.
5.1 Convergence
We begin by comparing the convergence of the methods discussed in this paper on several test problems. The results of this experiment are illustrated in Figure 2, and as expected, our augmented block-CG method outperforms the other methods, often by orders-of-magnitude. We also observe that Nyström PCG benefits from using $s>1$; i.e. from performing more than one passes over $\mathbf{A}$ when building the preconditioner.
As discussed in Section 1.1, throughout this paper we assume that matrix-loads (iterations) are the dominant cost. For reference, we have also show convergence as a function of matrix-vector products in Figure 9.
5.2 Regularization parameter
In Figure 3 we plot the error after a fixed number of matrix loads as a function of $\mu$. As expected, this plot indicates that our augmented block-CG outperforms the other methods for each value of $\mu$.
Perhaps more importantly, our block-CG method can efficiently compute the solution to Equation 1 for many values of $\mu$, without the need for additional matrix-load (see Theorem 3.7). This is in contrast to Nyström PCG which requires a new run for each value of $\mu$. For tasks such as ridge regression, this is of note.
Figure 3: Error after a fixed number of matrix-loads as a function of the regularization parameter $\mu$ for block-CG (Refer to caption), CG (Refer to caption), and Nyström PCG with $s=1$ (Refer to caption) and $s=3$ (Refer to caption).

### v2

@@ 3.2 Computational costs
We emphasize that after the block Lanczos algorithm has been run, the cost to compute the block-CG iterates for multiple values of $\mu$ is relatively small. This is in contrast to Nyström PCG, which is is not particularly well suited for solving Eq. 1 for multiple values of $\mu$; while the Nyström approximation can be reused, PCG must be re-run and new products with $\mathbf{A}$ computed.
Notice that the costs (besides matrix-loads) scale with the block-size $m$. We again remind the reader that, while Theorem 3.4 guarantees augmented block-CG outperforms Nyström preconditioning in terms of matrix-loads, but our theory provides no guarantees for other cost matrices like matrix-products or floating point operations. Users should carefully consider the costs and benefits associated with the various algorithms in the context of their own computing environments.
Remark 3.9.
By avoiding orthogonalizing and storing the whole Krylov basis, it is in fact possible to reduce the floating point cost to $O(dmtL+m^{3}tL)$ and the storage to $O(dm+dL)$, where $L$ is the number of values of $\mu$ at which one wishes to evaluate $\mathsf{bcg}_{t}^{(1)}(\mu)$ [48]. However, there are several caveats to such an implementation. First, such methods require knowing the values of $\mu$ at which one wishes to evaluate $\mathsf{bcg}_{t}^{(1)}(\mu)$ ahead of time. Second, the finite-precision behavior of such methods can differ greatly from the exact arithmetic behavior. We explore the impacts of finite precision arithmetic Appendix A, where we perform some numerical experiments that indicate that full reorthogonalization may be unnecessary in some situations.
For methods based on the standard Lanczos algorithm, there is quite a bit of theory which guarantees that such methods can still work, even without orthogonalization [49, 50, 19, 28, 46, 7]. Unfortunately, much less theory is known about the block-Lanczos algorithm in finite precision arithmetic [67]. Without reorthogonalization, we observe that in many cases the block-CG method fails to converge at all, while in other cases it does converge for some time; see Appendix A. Understanding this behavior is well beyond the scope of the current work.
4 Probabilistic bounds
In this section we prove probabilistic bounds for augmented block-CG. In light of Theorem 3.4, our strategy for deriving bounds for block-CG is simply to derive bounds for Nyström PCG. To do this, we follow the approach developed in [23] and use a deterministic bound for the condition number of the Nyström preconditioned system. The following is a minor generalization of Proposition 5.3 in [23] to allow arbitrary $\theta$. The proof is contained in Section C.1.
Proposition 4.1.
@@ 4 Probabilistic bounds
Let $\mathbf{P}_{\mu}$ be the Nyström preconditioner Eq. 9 corresponding to the Nyström approximation $\mathbf{A}\langle\mathbf{K}\rangle$ for any $\mathbf{K}$ and shift parameter $\theta\geq 0$. Then
@@
$\kappa(\mathbf{P}_{\mu}^{-1/2}\mathbf{A}_{\mu}\mathbf{P}_{\mu}^{-1/2})\leq\big{(}\theta+\mu+\|\mathbf{A}-\mathbf{A}\langle\mathbf{K}\rangle\|\big{)}\left(\frac{1}{\theta+\mu}+\frac{1}{\lambda_{d}+\mu}\right).$
In particular, Proposition 4.1 implies that if $\|\mathbf{A}-\mathbf{A}\langle\mathbf{K}\rangle\|\approx\lambda_{r+1}$ and $\theta\in[\lambda_{d},\lambda_{r+1}]$, then $\kappa(\mathbf{P}_{\mu}^{-1/2}\mathbf{A}_{\mu}\mathbf{P}_{\mu}^{-1/2})\approx\kappa_{r+1}(\mu)$.
Low-rank approximation is one of the most studied problems in randomized numerical linear algebra, and many bounds have been developed. In the remainder of this section, we employ these bounds to obtain guarantees for augmented block-CG.
Our first result, which we prove in Section 4.1, is a “numerical analysis” style bound based on the error gurantees for Nyström low-rank approximation in [61].
@@
$s\geq\min\left\{2+\frac{\log(d)}{2},{\frac{3}{2}}+\frac{1}{4}\log\left(4+\frac{4r}{p-1}\sum_{i>r}\frac{\lambda_{i}^{2}}{\lambda_{r+1}^{2}}\right)\right\}.$
@@
$\Bigg{\{}\forall\mu\geq 0:\frac{\|\mathbf{A}_{\mu}^{-1}\mathbf{b}-\mathsf{bcg}_{t}^{(1)}(\mu)\|_{\mathbf{A}_{\mu}}}{\|\mathbf{A}_{\mu}^{-1}\mathbf{b}\|_{\mathbf{A}_{\mu}}}\leq 2\exp\left(-\frac{t-s}{3\sqrt{\kappa_{r+1}(\mu)}}\right)\Bigg{\}}.$
This implies that, after a small burn-in period of at most $O(\log(d))$ iterations, we have exponential convergence at a rate roughly $1/\sqrt{\kappa_{r+1}(\mu)}$.
Note that $\mathbf{A}\langle\mathbf{K}_{s}\rangle$ as defined in Eq. 12 can have rank as large as $\ell s$, so we might hope that roughly $s=r/\ell$ iterations suffice. Recent work [44, 8] implies this is more-or-less the case. In Section 4.2, we use this result to prove our second result, a “theoretical computer science” type bound.
Theorem 4.3.
Let $\mathbf{\Omega}\in\mathbb{R}^{\ell}$ be a Gaussian matrix and $r\geq 0$. Then, for some
$s=O\left(\frac{r}{\ell}\log(\Delta)+\log\left(\frac{d}{\delta}\right)\right),\quad\Delta:=\min_{i=1,\ldots,\ell\lceil r/\ell\rceil-1}\frac{\lambda_{i}-\lambda_{i+1}}{\lambda_{1}},$
the block-CG iterate $\mathsf{bcg}_{t}^{(1)}(\mu)$ satisfies, with probability at least $1-\delta$,
$\Bigg{\{}\forall\mu\geq 0:\frac{\|\mathbf{A}_{\mu}^{-1}\mathbf{b}-\mathsf{bcg}_{t}^{(1)}(\mu)\|_{\mathbf{A}_{\mu}}}{\|\mathbf{A}_{\mu}^{-1}\mathbf{b}\|_{\mathbf{A}_{\mu}}}\leq 2\exp\left(-\frac{t-s}{3\sqrt{\kappa_{r+1}(\mu)}}\right)\Bigg{\}}.$
In contrast to Theorem 4.2, which requires $\ell\geq r+2$, Theorem 4.3 allows any choice of $\ell$. The bound reveals that we get exponential convergence at a rate roughly $1/\sqrt{\kappa_{r+1}(\mu)}$ after roughly $s\approx r/\ell$ matrix-loads (and hence $r\approx\ell s$ matrix-vector products). Note that the logarithmic dependence on eigenvalue gaps $\Delta$ is generally considered to be mild due to “smoothed-analysis” type effects of finite-precision arithmetic [44, 8].
Remark 4.4.
The fact that Theorems 4.2 and 4.3 guarantee an accurate result for all $\mu\geq 0$ with high probability (as opposed to a single value of $\mu$) will be important in our applications. In particular, it allows guarantees for computing the whole ridge-regression regularization path and will be necessary in the analysis of our algorithm for sampling Gaussian vectors.
Remark 4.5.
Bounds based on condition numbers (such as Theorem 4.2) are often pessimistic in practice. To determine how many iterations to run, it is more common to use some sort of a posteriori error estimate. For example, observing the residual $\|\mathbf{b}-\mathbf{A}_{\mu}\mathsf{bcg}_{t}^{(1)}(\mu)\|$ gives some indication of the quality of the solution. More advanced techniques can also be used [43].
4.1 Explicit bound
In this section, we prove Theorem 4.2. We do not attempt to optimize constants, opting instead for simple arguments and clean theorem statements.
We begin by recalling an error guarantee for Nyström low-rank approximation, that compares the Nyström error $\|\mathbf{A}-\mathbf{A}\langle\mathbf{K}_{s}\rangle\|$ to the error $\|\mathbf{A}-[\![\mathbf{A}]\!]_{r}\|=\lambda_{r+1}$ of the best possible rank-$r$ approximation to $\mathbf{A}$.
Theorem 4.6 (Theorem 9.1 in [61]).
Suppose $\mathbf{\Omega}\in\mathbb{R}^{d\times(r+p)}$ is a random Gaussian matrix and define $\mathbf{K}_{s}:=[\mathbf{\Omega}~\mathbf{A}\mathbf{\Omega}~\cdots~\mathbf{A}^{s-1}\mathbf{\Omega}]$. Then, if $p\geq 2$,
@@ 4.1 Explicit bound
$\log\Biggl{(}\frac{\operatorname{\mathbb{E}}\|\mathbf{A}-\mathbf{A}\langle\mathbf{K}_{s}\rangle\|^{2}}{\lambda_{r+1}^{2}}\Bigg{)}\leq\frac{1}{8(s-\frac{3}{2})^{2}}\log\Biggl{(}4+\frac{4r}{p-1}\sum_{i>r}\frac{\lambda_{i}^{2}}{\lambda_{r+1}^{2}}\Bigg{)}^{2}.$
To prove our main convergence guarantee for our augmented block-CG, we derive a new bound for Nyström preconditioning that may be of independent interest.
Theorem 4.7.
Suppose $\mathbf{\Omega}\in\mathbb{R}^{d\times(r+p)}$ is a random Gaussian matrix and define $\mathbf{K}_{s}:=[\mathbf{\Omega}~\mathbf{A}\mathbf{\Omega}~\cdots~\mathbf{A}^{s-1}\mathbf{\Omega}]$. Let $\mathbf{P}_{\mu}$ be the Nyström preconditioner Eq. 9 corresponding to the Nyström approximation $\mathbf{A}\langle\mathbf{K}_{s}\rangle$ for any shift parameter $\theta\in[\lambda_{d},\lambda_{r+1}]$. Suppose $p\geq 2$ and
$s\geq\min\left\{2+\frac{\log(d)}{2},\frac{3}{2}+\frac{1}{4}\log\left(4+\frac{4r}{p-1}\sum_{i>r}\frac{\lambda_{i}^{2}}{\lambda_{r+1}^{2}}\right)\right\}.$
Then, with probability at least $99/100$,
$\bigg{\{}\forall\mu\geq 0:\kappa(\mathbf{P}_{\mu}^{-1/2}\mathbf{A}_{\mu}\mathbf{P}_{\mu}^{-1/2})\leq 28\kappa_{r+1}(\mu)\bigg{\}}.$
Note that our bound for Nyström PCG applies even when $\mu=0$, whereas the bounds in [23] require $\mu$ to be sufficiently large relative to the sketching dimension $r+p$ and consider only $s=1$.
Proof 4.8.
The proof is a simple consequence of Theorems 4.6 and 4.1. We begin by bounding $\|\mathbf{A}-\mathbf{A}\langle\mathbf{K}\rangle\|$. We first note that the minimum is always attained by the second term. Indeed, since $r\leq d$ and $\lambda_{i}/\lambda_{r+1}\leq 1$ for $i>r$, by setting $p=2$, we see $4r/(p-1)\sum_{i>r}\lambda_{i}^{2}/\lambda_{r+1}^{2}\leq 4d^{2}$. Now, using properties of the logarithm, $\log(4+4d^{2})=\log(1+d^{2})+\log(4)\leq\log(d^{2})+\log(2)+\log(4)=\log(8)+2\log(d)$. The claim then follows since $3/2+\log(8)/4\leq 2$.
With the condition on $s$, Theorem 4.6 guarantees that
(25) $\operatorname{\mathbb{E}}\Bigl{[}\|\mathbf{A}-\mathbf{A}\langle\mathbf{K}\rangle\|^{2}\Bigr{]}\leq\mathrm{e}^{1/2}\cdot\lambda_{r+1}^{2}.$
Applying Markov’s inequality, we therefore have that
(26) $\operatorname{\mathbb{P}}\Bigl{[}\|\mathbf{A}-\mathbf{A}\langle\mathbf{K}\rangle\|^{2}\geq 100\cdot\mathrm{e}^{1/2}\lambda_{r+1}^{2}\Bigr{]}\leq\frac{1}{100}.$
Condition on the event that $\{\|\mathbf{A}-\mathbf{A}\langle\mathbf{K}\rangle\|\leq 10\mathrm{e}^{1/4}\lambda_{r+1}\}$. Since $\theta\leq\lambda_{r+1}$ and $\mu\geq 0$,
(27) $\theta+\mu+\|\mathbf{E}\|\leq(1+10\mathrm{e}^{1/4})(\lambda_{r+1}+\mu)\leq 14(\lambda_{r+1}+\mu).$
Next, since $\theta\geq\lambda_{d}$,
(28) $\left(\frac{1}{\theta+\mu}+\frac{1}{\lambda_{d}+\mu}\right)\leq\frac{2}{\lambda_{d}+\mu}.$
The result follows by combining the above equations.
The proof of Theorem 4.2 is now straightforward. To get a high probability bound, we use a simple boosting “trick”. We expect more refined results can be obtained directly.
Proof 4.9 (Proof of Theorem 4.2).
(29) $\Big{\{}\forall\mu\geq 0:\kappa(\mathbf{P}_{\mu}^{-1/2}\mathbf{A}_{\mu}\mathbf{P}_{\mu}^{-1/2})\leq 28\kappa_{r+1}(\mu)\Big{\}}.$
When the event in Eq. 29 holds, Corollary 2.3 guarantees that,
(30) $\Bigg{\{}\forall\mu\geq 0:\frac{\|\mathbf{A}^{-1}\mathbf{b}-\mathsf{pcg}_{t}(\mu)\|_{\mathbf{A}_{\mu}}}{\|\mathbf{A}^{-1}\mathbf{b}\|_{\mathbf{A}_{\mu}}}\leq 2\exp\left(-\frac{2t}{\sqrt{28\kappa_{r+1}(\mu)}}\right)\Bigg{\}}.$
The result follows from Theorem 3.4 and that $\sqrt{28}/2<3$.
We now analyze the case for arbitrary $q\geq 1$. Partition $\mathbf{\Omega}=[\mathbf{\Omega}_{1}~\cdots~\mathbf{\Omega}_{q}]\in\mathbb{R}^{d\times(r+2)q}$. Then, for $i=1,\ldots,q$, $\mathbf{\Omega}_{i}\in\mathbb{R}^{d\times(r+2)}$ are independent Gaussian matrices. By our analysis of the $q=1$ case, block-CG with starting block $\mathbf{b}~\mathbf{\Omega}_{i}]$ fails to reach the specified accuracy within $t$ iterations with probability at most $1/100$. The probability that all of these $q$ (independent) instances fail to reach this accuracy is therefore at most $(1/100)^{q}\leq\delta$. Finally, since
(31) $\mathcal{K}_{t}(\mathbf{A},[\mathbf{b}~\mathbf{\Omega}_{i}])\subseteq\mathcal{K}_{t}(\mathbf{A},\mathbf{\Omega})$
block-CG with starting block $\mathbf{b}~\mathbf{\Omega}]$ performs no worse than block-CG with starting block $\mathbf{b}~\mathbf{\Omega}_{i}]$ (for any $i$), and hence fails to reach the stated accuracy within $t$ iterations with probability at most $\delta$.
4.2 Improved matrix-vector product complexity
In this section, we prove Theorem 4.3. The general approach is identical to the proof of Theorem 4.2, but we use a different bound for Nyström low-rank approximation.
Theorem 4.10 (Theorem 1.3 in [8]).
Let $\mathbf{\Omega}\in\mathbb{R}^{d\times\ell}$ be a Gaussian matrix. Then, for some
$s=O\left(\frac{r}{\ell\sqrt{\varepsilon}}\log(\Delta)+\log\left(\frac{d}{\delta\varepsilon}\right)\right),\quad\Delta:=\min_{i=1,\ldots,\ell\lceil r/\ell\rceil-1}\frac{\lambda_{i}-\lambda_{i+1}}{\lambda_{1}},$
with probability at least $1-\delta$, there is a matrix $\mathbf{Q}\in\mathbb{R}^{d\times r}$ with orthonormal columns and $\operatorname{range}(\mathbf{Q})\subseteq\mathcal{K}_{s}(\mathbf{A},\mathbf{\Omega})$ such that
$\|\mathbf{A}-\mathbf{Q}\mathbf{Q}^{\mathsf{T}}\mathbf{A}\|<(1+\varepsilon)\lambda_{r+1}.$
Note that Theorem 4.10 is not for Nyström low-rank approximation. However, Nyström low-rank approximation is better than projection-based low-rank approximation [61], which allows us to prove Theorem 4.3.
Proof 4.11 (Proof of Theorem 4.3).
We set $\varepsilon=10\mathrm{e}^{1/4}-1$. Since $\operatorname{range}(\mathbf{Q})\subseteq\operatorname{range}(\mathbf{K}_{s})$, the standard monotonicity property of the Nyström approximation implies $\|\mathbf{A}-\mathbf{A}\langle\mathbf{K}_{s}\rangle\|\leq\|\mathbf{A}-\mathbf{A}\langle\mathbf{Q}\rangle\|$. Next, [61, Lemma 5.2] asserts that for any matrix $\mathbf{Q}$ with orthonormal columns, $\|\mathbf{A}-\mathbf{A}\langle\mathbf{Q}\rangle\|\leq\|\mathbf{A}-\mathbf{Q}\mathbf{Q}^{\mathsf{T}}\mathbf{A}\|$. Thus, by Theorem 4.10 (with $\varepsilon$ set to an appropriate constant), the given condition on $s$ ensures that, with probability at least $1-\delta$, $\|\mathbf{A}-\mathbf{A}\langle\mathbf{K}_{s}\rangle\|<10\mathrm{e}^{1/4}\lambda_{r+1}$. The proof then follows similarly to the proof of Theorem 4.2.
5 Discussion and comparison with past work
The block-CG algorithm is widely used to solve linear systems with multiple right hand sides, and in such settings working over the block Krylov subspace is often advantageous compared to working over the individual Krylov subspaces [48, 41, 22, 4]. Block-CG is also used to solve single linear systems. For instance, so-called enlarged KSMs split the right hand size vector $\mathbf{b}$ into multiple vectors using a domain decomposition approach, and then apply block-CG to this collection of vectors [30, 31, 32]. Augmenting a KSM with random vectors was shown to be beneficial in numerical experiments appearing in concurrent work [66, Appendix B].
While we are unaware of any bounds similar to Theorem 4.3, bounds similar to Theorem 4.2 are known. In particular, [48, §4] derives a convergence bound for block-CG terms of $\sqrt{\lambda_{1}/\lambda_{d-m+1}}$; i.e. the condition number of $\mathbf{A}$ with the bottom $m$ eigenvalues removed. [^5] Note that the ordering of the eigenvalues in [48] is reversed from the present paper. At a high level, [48] obtains this bound by showing that the block Krylov subspace contains vectors which can approximately annihilate a subset of the eigenvalues of $\mathbf{A}$. Thus, in principle, other sets of eigenvalues (besides the bottom ones) can be eliminated. The bounds of [48] work for any starting block (but as a result, are more technical, depending on how “good” the starting block is). It would not be surprising such quantities can be bounded when the starting block is Gaussian.
KSMs are widely used for both low-rank approximation and problems involving matrix-functions, but their behavior on these problems is fundamentally different. As such, it would be interesting to try to avoid relying on a black-box reduction to low-rank approximation. For example, while low-rank approximation is complicated by the presence of repeated eigenvalues, solving linear systems is not. The approach of [48] is closely related to the approaches used in [61, 45] to prove bounds for low-rank approximation, so it is conceivable that, by taking the approach of [48], but using modern tools from randomized numerical linear algebra, one may be able to directly derive better bounds for randomized versions of block-CG.
More broadly, there are many randomized algorithms for linear systems and regression. For instance, a recent line of work shows that certain stochastic iterative methods enjoy an implicit preconditioning-like effect [14, 39, 17, 18, 16], with convergence independent of the top eigenvalues. These algorithms have runtime guarantees better than KSMs in some settings [16]. In addition, for tall least squares problems, the sketch-and-precondition framework for building a preconditioner for KSMs is extremely effective [54, 40], resulting in algorithms with optimal theoretical complexity [13, 11]. A number of recent works in randomized numerical linear algebra [10, 61, 44, 8] have made use of various nestedness properties of block-KSMs similar in flavor to Theorem 3.2 in order to obtain more efficient algorithms.